How do I get the answer to Calculate the following binomial probability by either using one of the binomial probability tables, software, or a calculator using the formula below. Round your answer to 3 decimal places.

P(x | n, p) =

n!
(nx)! x!

· px · qnx    where    q = 1 − p

P(x < 8, n = 9, p = 0.9)

=
Answer #1

Binomial Formula: P(X) = nCx px qn-x

n = 9

p = 0.9

q = 1 - p = 0.1

P(X < 8) = 1 - P(X \geq 8)

= 1 - [P(8) + P(9)]

= 1 - [(9C8 x 0.98 x 0.1) + (9C9 x 0.99)]

= 1 - [(9!/(8!x1!)) x 0.98 x 0.1) + ((9!/9!) x 0.99)]

= 1 - [(9 x 0.98 x 0.1) + (1 x 0.99)]

= 1 - (0.3874 + 0.3874)

= 0.225